Jacob has a box of bills. There are five more $10 bills than $5 bills, but two less $2 bills than $5 bills. He saved $148. How many of each bill did he save?
Answer by ptaylor(2198)

Let x=number of $5 bills; 5x=value of $5 bills
Then x+5=number of $10 bills; 10(x+5)=value of $10 bills
And x-2=number of $2 bills; 2(x-2)=value of $2 bills
Now we are told that:
5x+10(x+5)+2(x-2)=148 get rid of parens (distributive law)
5x+10x+50+2x-4=148 collect like terms
17x+46=148 subtract 46 from both sides
17x+46-46=148-46 collect like terms again
17x=102 divide both sides by 17
x=6----------------------------------number of $5 bills
x+5=6+5=11--------------------------number of $10 bills
x-2=6-2=4-------------------------number of $2 bills
CK
6*5+11*10+4*2=148
30+110+8=148
148=148

Hope this helps---ptaylor